OpenStudy (anonymous):

Question about n^x and the derivative. So I know that if f(x)=n^x, then f'(x)=n^x*ln(n). I don't really understand WHY though? I know the PROOFS, but I don't have a physical understanding as to WHY this is true. So if you have a number n raised to the power of x, the slope of the tangent line at the point n^x is y value of n^x, multiplied by the exponent required for base e to equal n? That's what I'm not understanding.

OpenStudy (amistre64):

let y = n^x we can take the log of each side log(y) = log(n^x) using log properties we get log(y) = x log(n) taking the derivative of each side we get y'/y = log(n) solve for y' y' = y log(n) but we know from the start that y = n^x, therefore y' = n^x log(n)

OpenStudy (anonymous):

really depends on your definition of \(f(x)=b^x\) one is \(b^x=e^{x\ln(b)}\) so for example \(f(x)=2^x=e^{x\ln(2)}\) then take the derivative and you will see it right away @amistre64 method also explains it

OpenStudy (amistre64):

if n=e that reduces us to the (e^x)' = e^x log(e); or simply (e^x)' = e^x

OpenStudy (anonymous):

the definition \[b^x=e^{x\ln(b)}\] makes sense out of things like \(2^{\pi}\) for example, it becomes \[e^{\pi\ln(2)}\] it also allows you to make sense out of functions that have a variable in the base and the exponent. for example \[x^{\sin(x)}=e^{\sin(x)\ln(x)}\]

OpenStudy (amistre64):

real analysis defines:\[log(x)=\int_{1}^{\infty}\frac1xdx\] earlier math classes use a variety of log functions differentiated by bases ... but since they can all be configured in log(x) form, they simply define a "log" function. such that log(e) = 1

OpenStudy (anonymous):

I understand the algebra behind it, I've seen many proofs. I'm trying to understand how the slope is related to "the exponent that e must be raised by to equal n". I understand how n^x can affect a slope, that's very obvious. But why is ln(n) altering the slope? What's that relation?

OpenStudy (amistre64):

most likely your trying to consider how the chain rule works. consider a box full of gears such that you can turn the first gear at a certain rate and it effects the last gears output rate. a transmission in a car if you will.

OpenStudy (amistre64):

e^u -> u' e^u ; but then we have to ask ourselves, what is effecting u? let u = x ln(n) ; u' = ln(n) for suitable constant n